# On estimation and order selection for multivariate extremes via clustering

### v3

@@ 1 Introduction
Recently, as attempts that can also be viewed as providing a parsimonious summary of the angular distribution of multivariate extremes, several authors considered applying clustering algorithms over the sphere on which the spectral measure resides. Einmahl et al. [7] and Janßen and Wan [23] applied the spherical $k$-means algorithm based on cosine dissimilarity [6] and addressed its relation to the estimation of max-linear factor models. Fomichov and Ivanovs [10] proposed the spherical $k$-principal-component ($k$-pc) algorithm which is based on a modified cosine dissimilarity, and discussed its superiority in terms of detecting the concentration of the spectral measure on lower-dimensional faces. Medina et al. [26] considered applying the spectral clustering algorithm [27] to the $k$-nearest neighbor graph constructed from the angular part of the extreme samples, and related it to sum-linear factor models.
As readily observed in the aforementioned works, there is a natural connection between a discrete spectral measure and spherical clustering: Each atom in the spectral measure can be viewed as a cluster center (prototype), and the angular part of the extreme samples form clusters around these atoms. In fact, this intuition has been rigorously explored by [23] and [26], where consistent recovery of the spectral measure based on their clustering algorithms was established (the consistency result of [23] also applies to the $k$-pc clustering of [10]). Since models such as the heavy-tailed max-linear and sum-linear factor models are essentially characterized by the spectral measure, the consistent estimation of spectral measure can be, in principle, converted to the consistent estimation of parameters of the factor models.
So far, in all the theoretical analysis of the works linking a discrete spectral measure to a clustering algorithm, the number of atoms, or equivalently speaking, the number of clusters, is assumed to be known. We refer to this number as the order, since it also relates to the order of the factor models mentioned above. In [23, 10, 26], ad hoc methods such as elbow plot and scree plot were used to guide the selection of the order in their real data analysis. These ad hoc methods are based solely on human visuals to locate the vaguely defined “elbow” point, and lack theoretical justification.
@@ 2.1 Multivariate extreme value theory
$\lim_{x\rightarrow\infty}x^{\alpha}\Pr(X_{1}>x)=\ldots=\lim_{x\rightarrow\infty}x^{\alpha}\Pr(X_{d}>x)=1,$ (1)
@@
where $[\mathbf{0},\mathbf{x}]=[0,x_{1}]\times\ldots\times[0,x_{d}]$, $x_{i}>0$, $i\in\left\{1,\ldots,d\right\}$. Moreover, the measure $\Lambda$ satisfies the homogeneity property $\Lambda(c\,\cdot)=c^{-\alpha}\Lambda(\cdot)$, $c>0$, and therefore admits a polar decomposition into a product of radial and angular parts. We shall follow the formulation in [1, Section 8.2.5], which allows the use of different norms for the radial and angular components. Suppose that $\|\cdot\|_{(r)}$ and $\|\cdot\|_{(s)}$ denote two arbitrary norms on $\mathbb{R}^{d}$. Slightly abusing the notation, we still use $\Lambda$ denote the push-forward measure of $\Lambda$ under the one-to-one mapping $\mathbb{E}_{d}\mapsto(0,\infty)\times\mathbb{S}_{+}^{d-1}$, $\mathbf{x}\mapsto(r,\mathbf{w}=(w_{1},\ldots,w_{d})):=\left(\|\mathbf{x}\|_{(r)},\mathbf{x}/\|\mathbf{x}\|_{(s)}\right)$, where
@@
$\int_{\mathbb{S}_{+}^{d-1}}\left(\frac{w_{1}}{\|\mathbf{w}\|_{(r)}}\right)^{\alpha}H(d\mathbf{w})=\ldots=\int_{\mathbb{S}_{+}^{d-1}}\left(\frac{w_{d}}{\|\mathbf{w}\|_{(r)}}\right)^{\alpha}H(d\mathbf{w})=\frac{1}{c_{(r)}}.$ (7)
In practice, commonly used norms are the $p$-norm $\|\mathbf{x}\|_{p}=\left(\sum_{j=1}^{d}|x_{j}|^{p}\right)^{1/p}$, $p\in(0,\infty)$, and the sup-norm $\|\mathbf{x}\|_{\infty}=\max\left(|x_{1}|,\ldots,|x_{d}|\right)$. In view of (2), the following weak convergence on $\mathbb{S}_{+}^{d-1}$ holds
@@ 2.2 Spherical clustering
A dissimilarity measure $D$ on $\mathbb{S}_{+}^{d-1}$ is a bivariate function $D:\mathbb{S}_{+}^{d-1}\times\mathbb{S}_{+}^{d-1}\rightarrow[0,1]$ that is continuous and satisfies the following properties: for $\mathbf{w}_{i}\in\mathbb{S}_{+}^{d-1}$, $i\in\left\{1,2\right\}$, (i) $D(\mathbf{w}_{1},\mathbf{w}_{2})=0$ if and only if $\mathbf{w}_{1}=\mathbf{w}_{2}$; (ii) $D(\mathbf{w}_{1},\mathbf{w}_{2})=D(\mathbf{w}_{2},\mathbf{w}_{1})$.
@@
To simplify the mathematical description of clustering of sample data, it is convenient to use the notion of multiset (e.g., [24]). Recall that a multiset $W$ on $\mathbb{S}_{+}^{d-1}$ is a set that allows repetition of its elements, whose support, denoted as $\operatorname{supp}{W}$, is a subset of $\mathbb{S}_{+}^{d-1}$ in the usual sense that eliminates repetitions in $W$. For instance, with two distinct points $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$ on $\mathbb{S}_{+}^{d-1}$, one can have $W=\{\mathbf{w}_{1},\mathbf{w}_{1},\mathbf{w}_{2}\}$ with $\operatorname{supp}{W}=\{\mathbf{w}_{1},\mathbf{w}_{2}\}$. A multiset $W$ can be characterized by the multiplicity function $m_{W}:\mathbb{S}_{+}^{d-1}\mapsto\{0,1,2,\ldots\}$, where $m_{W}(\mathbf{w})$ equals the number of repetitions of element $\mathbf{w}\in\mathbb{S}_{+}^{d-1}$ ($m_{W}(\mathbf{w})=0$ if $\mathbf{w}\notin\operatorname{supp}{W}$). A subset of $\mathbb{S}_{+}^{d-1}$ in the usual sense can be understood as a multiset with the multiplicity taking value either 00 or $1$, with the empty set corresponding to a multiplicity function that is identically 00. When the notation $\mathbf{w}\in W$ is used for a multiset $W$, it means that $\mathbf{w}$ is an element in $\operatorname{supp}{W}$. For multisets $W_{1},W_{2}$ with multiplicity functions $m_{1}$ and $m_{2}$ respectively, their union $W_{1}\cup W_{2}$ is given by the multiset characterized by the multiplicity function $m_{1}\vee m_{2}$, and their intersection $W_{1}\cap W_{2}$ is given by the multiset characterized by $m_{1}\wedge m_{2}$. The relation $W_{1}\subset W_{2}$ is understood as $m_{1}\leq m_{2}$. Furthermore, if $\operatorname{supp}{W}$ is a finite set, a summation $\sum_{\mathbf{w}\in W}f(\mathbf{w})$ for a suitable function $f$ is understood as $\sum_{\mathbf{w}\in\operatorname{supp}{W}}f(\mathbf{w})m_{W}(\mathbf{w})$. For example, the cardinality of $W$ is defined as
@@
The existence of $A_{k}^{*}$ is guaranteed by the continuity of $D$ and the compactness of $\mathbb{S}_{+}^{d-1}$, although it may not be unique. Notice that when $|\operatorname{supp}{W}|\geq k$, the infimum in (13) must be achieved with a distinct set of $\mathbf{a}_{i}^{*}$’s. Below when multisets $C_{1},\ldots,C_{k}$ with multiplicity functions $m_{1},\ldots,m_{k}$ are said to form a partition of a multiset $W$ with multiplicity function $m$, it means that $m=m_{1}+\ldots+m_{k}$, and $m_{i}\neq 0$ for all $i\in\{1,\cdots,k\}$.
@@
A $k$-clustering of a multiset $W$, $1\leq k\leq|W|$, with respect to the dissimilarity measure $D$ refers to a pair $(A_{k}^{*},\mathfrak{C}_{k})$. Here $A_{k}^{*}=\left\{\mathbf{a}_{1}^{*},\ldots,\mathbf{a}_{k}^{*}\right\}$ is as described above, and $\mathfrak{C}_{k}=\{C_{1},\ldots,C_{k}\}$ is a partition of $W$ into a collection of multisets $C_{i}$’s such that $D\left(\mathbf{w},A_{k}^{*}\right)=D\left(\mathbf{w},\mathbf{a}_{i}^{*}\right)$ for all $\mathbf{w}\in C_{i}$, $i\in\left\{1,\ldots,k\right\}$. We refer to $A_{k}^{*}$ as the set of centers and each $C_{i}$ as a cluster.
@@ 2.3 Spherical clustering for multivariate extremes
where $\mathbf{a}_{i}$’s are distinct points on $\mathbb{S}_{+}^{d-1}$, and $p_{i}>0$, $p_{1}+\ldots+p_{k}=1$. Let $W_{n}$ denote the extremal subsample as in (14), and $\left({A}_{k,n}=\left({\mathbf{a}}_{1,n}^{k},\ldots,{\mathbf{a}}_{k,n}^{k}\right),\mathfrak{C}_{k,n}=\left\{{C}_{1,n}^{k},\ldots,{C}_{k,n}^{k}\right\}\right)$ form a $k$-clustering of $W_{n}$ as defined in Definition 2 with respect to a dissimilarity measure $D$ defined in Definition 1. Let
@@
$\mathbf{a}_{\pi_{n}(i),n}^{k}\rightarrow\mathbf{a}_{i},\ \text{ and }\ p_{\pi_{n}(i),n}^{k}\rightarrow p_{i},\quad i\in\{1,\ldots,k\},$
@@
Proof.
The convergence of $\mathbf{a}_{\pi_{n}(i),n}^{k}$ follows from [23, Theorem 3.1] (stated as convergence in Hausdorff distance between ${A}_{k,n}$ and $\{\mathbf{a}_{1},\ldots,\mathbf{a}_{k}\}$) and Proposition 1 above; see also the discussion in [23, Section 4]). It remains to show the convergence of $p_{\pi_{n}(i),n}^{k}$, $i\in\{1,\ldots,k\}$. Set
@@
Comparing [23, Proposition 3.3] with Proposition 1 and Corollary 1 here, we have chosen to work directly under the marginal standardization assumption in (1) and not to treat the empirical marginal transformations as in [23, Eq. (3.5)] for simplicity. Nevertheless, the consistency result of the order selection below (Theorem 1) can be extended to the setup of [23] based on the results there.
@@ 3.1 The method
$a(\mathbf{w})=D\left(\mathbf{w},A_{k}^{*}\right),\quad\text{and}\quad b(\mathbf{w})=\bigvee_{i=1}^{k}D\left(\mathbf{w},A_{k}^{*}\setminus\left\{\mathbf{a}_{i}^{*}\right\}\right),$
@@
A well-clustered dataset is expected to have small $a(\mathbf{w})$ values relative to $b(\mathbf{w})$ across the majority of $\mathbf{w}$ points. Hence, one often uses $\bar{S}$ to guide the selection of the number of clusters, that is, to choose $k$ which maximizes $\bar{S}$. However, when experimenting applying the ASW to multivariate extremes with a discrete spectral measure as described in Section 2.3, the performance is unsatisfactory: it tends to respond insensitively when the number of clusters exceeds the true $k$, i.e., the number of atoms of the spectral measure; see, for example, the curve corresponding to $t=0$ in Figure 1. In particular, we observe two behaviors of ASW that lead to the issue: 1) it tends to treat a tiny fraction of isolated points as a cluster; 2) it sometimes splits a single cluster center into multiple centers that are close to each other.
@@
Suppose $\mathbf{X}$ satisfying (1) and (2) has a discrete spectral measure of the form $H=\sum_{i=1}^{k}p_{i}\delta_{\mathbf{a}_{i}}$, $k\in\mathbb{Z}_{+}$ where $\mathbf{a}_{i}$’s are distinct points on $\mathbb{S}_{+}^{d-1}$, and $p_{i}>0$, $p_{1}+\ldots+p_{k}=1$. Let $W_{n}$ denote the extremal subsample as in (14) with $\ell_{n}\rightarrow\infty$, $\ell_{n}/n\rightarrow 0$, and $\left(A_{m,n},\mathfrak{C}_{m,n}\right)$, $m\in\mathbb{Z}_{+}$, form an $m$-clustering of $W_{n}$ as defined in Definition 2 with respect to a dissimilarity measure $D$ defined in Definition 1. Let $r_{A}$ be defined as in (20) and define
@@
In practice, we suggest plotting the penalized ASW $S_{t}$ as a function of $m=1,2,\ldots$, for a range of small $t$ values. The idea is to start with $t$ near 00, gradually increase it, and see how the penalized ASW curve responds. If some obvious spurious clusters (i.e., those with tiny size or centers that are too close together) are present, the curve tends to respond sensitively and bends at the appropriate order. We then identify the turning point $m$ as the choice of the order $k$. As a quick illustration, we follow a simulation setup of $(d=6,k=6)$ described in Section 6.1 below to simulate a max-linear factor model (Section 5.1). See Figure 1. Increasing $t$ to a very large value is not informative and is not recommended in practice. On the other hand, it would be desirable to develop a data-driven method for choosing $t$, which we leave for a future work to explore.

### v4

@@ 1 Introduction
Recently, as attempts that can also be viewed as providing a parsimonious summary of the angular distribution of multivariate extremes, several authors considered applying clustering algorithms over the sphere on which the spectral measure resides. Einmahl et al. [6] and Janßen and Wan [22] applied the spherical $k$-means algorithm based on cosine dissimilarity [5] and addressed its relation to the estimation of max-linear factor models. Fomichov and Ivanovs [9] proposed the spherical $k$-principal-component ($k$-pc) algorithm which is based on a modified cosine dissimilarity, and discussed its superiority in terms of detecting the concentration of the spectral measure on lower-dimensional faces. medina2024spectral considered applying the spectral clustering algorithm [25] to the $k$-nearest neighbor graph constructed from the angular part of the extreme samples, and related it to sum-linear factor models.
As readily observed in the aforementioned works, there is a natural connection between a discrete spectral measure and spherical clustering: Each atom in the spectral measure can be viewed as a cluster center (prototype), and the angular part of the extreme samples form clusters around these atoms. In fact, this intuition has been rigorously explored by [22, medina2024spectral], where consistent recovery of the spectral measure based on their clustering algorithms was established (the consistency result of [22] also applies to the $k$-pc clustering of [9]). Since models such as the heavy-tailed max-linear and sum-linear factor models are essentially characterized by the spectral measure, the consistent estimation of spectral measure can be, in principle, converted to the consistent estimation of parameters of the factor models.
So far, in all the theoretical analysis of the works linking a discrete spectral measure to a clustering algorithm, the number of atoms, or equivalently speaking, the number of clusters, is assumed to be known. We refer to this number as the order, since it also relates to the order of the factor models mentioned above. In [22, 9, medina2024spectral], ad hoc methods such as elbow plot and scree plot were used to guide the selection of the order in their real data analysis. These ad hoc methods are based solely on human visuals to locate the vaguely defined “elbow” point, and lack theoretical justification.
@@ 2.1 Multivariate extreme value theory
$\lim_{x\rightarrow\infty}x^{\alpha}\Pr(X_{1}>x)=\cdots=\lim_{x\rightarrow\infty}x^{\alpha}\Pr(X_{d}>x)=1,$ (1)
@@
where $[\mathbf{0},\mathbf{x}]=[0,x_{1}]\times\cdots\times[0,x_{d}]$, $x_{i}>0$, $i\in\left\{1,\ldots,d\right\}$. Moreover, the measure $\Lambda$ satisfies the homogeneity property $\Lambda(c\,\cdot)=c^{-\alpha}\Lambda(\cdot)$, $c>0$, and therefore admits a polar decomposition into a product of radial and angular parts. We shall follow the formulation in [1, Section 8.2.5], which allows the use of different norms for the radial and angular components. Suppose that $\|\cdot\|_{(r)}$ and $\|\cdot\|_{(s)}$ denote two arbitrary norms on $\mathbb{R}^{d}$. Slightly abusing the notation, we still use $\Lambda$ denote the push-forward measure of $\Lambda$ under the one-to-one mapping $\mathbb{E}_{d}\mapsto(0,\infty)\times\mathbb{S}_{+}^{d-1}$, $\mathbf{x}\mapsto(r,\mathbf{w}=(w_{1},\ldots,w_{d})):=\left(\|\mathbf{x}\|_{(r)},\mathbf{x}/\|\mathbf{x}\|_{(s)}\right)$, where
@@
$\int_{\mathbb{S}_{+}^{d-1}}\left(\frac{w_{1}}{\|\mathbf{w}\|_{(r)}}\right)^{\alpha}H(d\mathbf{w})=\cdots=\int_{\mathbb{S}_{+}^{d-1}}\left(\frac{w_{d}}{\|\mathbf{w}\|_{(r)}}\right)^{\alpha}H(d\mathbf{w})=\frac{1}{c_{(r)}}.$ (7)
In practice, commonly used norms are the $p$-norm $\|\mathbf{x}\|_{p}=\left(\sum_{j=1}^{d}|x_{j}|^{p}\right)^{1/p}$, $p\in(0,\infty)$, and the sup-norm $\|\mathbf{x}\|_{\infty}=\max\left(|x_{1}|,\ldots,|x_{d}|\right)$. In view of (2), the following weak convergence on $\mathbb{S}_{+}^{d-1}$ holds, as $u\rightarrow\infty$,
@@ 2.2 Spherical clustering
A dissimilarity measure $D$ on $\mathbb{S}_{+}^{d-1}$ is a bivariate function $D:\mathbb{S}_{+}^{d-1}\times\mathbb{S}_{+}^{d-1}\rightarrow[0,1]$ that is continuous and satisfies the following properties: (i) $D(\mathbf{w}_{1},\mathbf{w}_{2})=0$ if and only if $\mathbf{w}_{1}=\mathbf{w}_{2}$, and (ii) $D(\mathbf{w}_{1},\mathbf{w}_{2})=D(\mathbf{w}_{2},\mathbf{w}_{1})$, where $\mathbf{w}_{i}\in\mathbb{S}_{+}^{d-1}$, $i\in\left\{1,2\right\}$.
@@
To simplify the mathematical description of clustering of sample data, it is convenient to use the notion of multiset (e.g., [23]). Recall that a multiset $W$ on $\mathbb{S}_{+}^{d-1}$ is a set that allows repetition of its elements, whose support, denoted as $\operatorname{supp}{W}$, is a subset of $\mathbb{S}_{+}^{d-1}$ in the usual sense that eliminates repetitions in $W$. For instance, with two distinct points $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$ on $\mathbb{S}_{+}^{d-1}$, one can have $W=\{\mathbf{w}_{1},\mathbf{w}_{1},\mathbf{w}_{2}\}$ with $\operatorname{supp}{W}=\{\mathbf{w}_{1},\mathbf{w}_{2}\}$. A multiset $W$ can be characterized by the multiplicity function $m_{W}:\mathbb{S}_{+}^{d-1}\mapsto\{0,1,\ldots\}$, where $m_{W}(\mathbf{w})$ equals the number of repetitions of element $\mathbf{w}\in\mathbb{S}_{+}^{d-1}$ ($m_{W}(\mathbf{w})=0$ if $\mathbf{w}\notin\operatorname{supp}{W}$). A subset of $\mathbb{S}_{+}^{d-1}$ in the usual sense can be understood as a multiset with the multiplicity taking value either 00 or $1$, with the empty set corresponding to a multiplicity function that is identically 00. When the notation $\mathbf{w}\in W$ is used for a multiset $W$, it means that $\mathbf{w}$ is an element in $\operatorname{supp}{W}$. For multisets $W_{1},W_{2}$ with multiplicity functions $m_{1}$ and $m_{2}$ respectively, their union $W_{1}\cup W_{2}$ is given by the multiset characterized by the multiplicity function $m_{1}\vee m_{2}$, and their intersection $W_{1}\cap W_{2}$ is given by the multiset characterized by $m_{1}\wedge m_{2}$. The relation $W_{1}\subset W_{2}$ is understood as $m_{1}\leq m_{2}$. Furthermore, if $\operatorname{supp}{W}$ is a finite set, a summation $\sum_{\mathbf{w}\in W}f(\mathbf{w})$ for a suitable function $f$ is understood as $\sum_{\mathbf{w}\in\operatorname{supp}{W}}f(\mathbf{w})m_{W}(\mathbf{w})$. For example, the cardinality of $W$ is defined as
@@
The existence of $A_{k}^{*}$ is guaranteed by the continuity of $D$ and the compactness of $\mathbb{S}_{+}^{d-1}$, although it may not be unique. Notice that when $|\operatorname{supp}{W}|\geq k$, the infimum in (13) must be achieved with a distinct set of $\mathbf{a}_{i}^{*}$’s. Below when multisets $C_{1},\ldots,C_{k}$ with multiplicity functions $m_{1},\ldots,m_{k}$ are said to form a partition of a multiset $W$ with multiplicity function $m$, it means that $m=m_{1}+\cdots+m_{k}$, and $m_{i}\neq 0$ for all $i\in\{1,\ldots,k\}$.
@@
A $k$-clustering of a multiset $W$, $1\leq k\leq|W|$, with respect to the dissimilarity measure $D$ refers to a pair $(A_{k}^{*},\mathfrak{C}_{k})$. Here $A_{k}^{*}=\left\{\mathbf{a}_{1}^{*},\cdots,\mathbf{a}_{k}^{*}\right\}$ is as described above, and $\mathfrak{C}_{k}=\{C_{1},\ldots,C_{k}\}$ is a partition of $W$ into a collection of multisets $C_{i}$’s such that $D\left(\mathbf{w},A_{k}^{*}\right)=D\left(\mathbf{w},\mathbf{a}_{i}^{*}\right)$ for all $\mathbf{w}\in C_{i}$, $i\in\left\{1,\ldots,k\right\}$. We refer to $A_{k}^{*}$ as the set of centers and each $C_{i}$ as a cluster.
@@ 2.3 Spherical clustering for multivariate extremes
where $\mathbf{a}_{i}$’s are distinct points on $\mathbb{S}_{+}^{d-1}$, and $p_{i}>0$, $p_{1}+\cdots+p_{k}=1$. Let $W_{n}$ denote the extremal subsample as in (14), and $\left({A}_{k,n}=\left({\mathbf{a}}_{1,n}^{k},\ldots,{\mathbf{a}}_{k,n}^{k}\right),\mathfrak{C}_{k,n}=\left\{{C}_{1,n}^{k},\ldots,{C}_{k,n}^{k}\right\}\right)$ form a $k$-clustering of $W_{n}$ as defined in Definition 2 with respect to a dissimilarity measure $D$ defined in Definition 1. Let
@@
$\mathbf{a}_{\pi_{n}(i),n}^{k}\rightarrow\mathbf{a}_{i},\quad p_{\pi_{n}(i),n}^{k}\rightarrow p_{i},\quad i\in\{1,\ldots,k\},$
@@
Proof:.
The convergence of $\mathbf{a}_{\pi_{n}(i),n}^{k}$ follows from [22, Theorem 3.1] (stated as convergence in Hausdorff distance between ${A}_{k,n}$ and $\{\mathbf{a}_{1},\ldots,\mathbf{a}_{k}\}$) and Proposition 1 above; see also the discussion in [22, Section 4]). It remains to show the convergence of $p_{\pi_{n}(i),n}^{k}$, $i\in\{1,\ldots,k\}$. Set
@@
Comparing [22, Proposition 3.3] with Proposition 1 and Corollary 1 here, we have chosen to work directly under the marginal standardization assumption in (1) and not to treat the empirical marginal transformations as in [22, Eq. (3.5)] for simplicity. Nevertheless, the consistency result of the order selection below (Theorem 1) can be extended to the setup of [22] based on the results there.
@@ 3.1 The method
$a(\mathbf{w})=D\left(\mathbf{w},A_{k}^{*}\right),\quad b(\mathbf{w})=\bigvee_{i=1}^{k}D\left(\mathbf{w},A_{k}^{*}\setminus\left\{\mathbf{a}_{i}^{*}\right\}\right),$
@@
A well-clustered dataset is expected to have small $a(\mathbf{w})$ values relative to $b(\mathbf{w})$ across the majority of $\mathbf{w}$ points. Hence, one often uses $\bar{S}$ to guide the selection of the number of clusters, that is, to choose $k$ which maximizes $\bar{S}$. However, when experimenting applying the ASW to multivariate extremes with a discrete spectral measure as described in Section 2.3, the performance is unsatisfactory: it tends to respond insensitively when the number of clusters exceeds the true $k$, i.e., the number of atoms of the spectral measure; see, for example, the curve corresponding to $t=0$ in Fig. 1. In particular, we observe two behaviors of ASW that lead to the issue: 1) it tends to treat a tiny fraction of isolated points as a cluster; 2) it sometimes splits a single cluster center into multiple centers that are close to each other.
@@
Suppose $\mathbf{X}$ satisfying (1) and (2) has a discrete spectral measure of the form $H=\sum_{i=1}^{k}p_{i}\delta_{\mathbf{a}_{i}}$, $k\in\mathbb{Z}_{+}$ where $\mathbf{a}_{i}$’s are distinct points on $\mathbb{S}_{+}^{d-1}$, and $p_{i}>0$, $p_{1}+\cdots+p_{k}=1$. Let $W_{n}$ denote the extremal subsample as in (14) with $\ell_{n}\rightarrow\infty$, $\ell_{n}/n\rightarrow 0$, and $\left(A_{m,n},\mathfrak{C}_{m,n}\right)$, $m\in\mathbb{Z}_{+}$, form an $m$-clustering of $W_{n}$ as defined in Definition 2 with respect to a dissimilarity measure $D$ defined in Definition 1. Let $r_{A}$ be defined as in (20) and define
@@
In practice, we suggest plotting the penalized ASW $S_{t}$ as a function of $m\in\{1,2,\ldots\}$ for a range of small $t$ values. The idea is to start with $t$ near 00, gradually increase it, and see how the penalized ASW curve responds. If some obvious spurious clusters (i.e., those with tiny size or centers that are too close together) are present, the curve tends to respond sensitively and bends at the appropriate order. We then identify the turning point $m$ as the choice of the order $k$. As a quick illustration, we follow a simulation setup of $(d=6,k=6)$ described in Section 6.1 below to simulate a max-linear factor model (Section 5.1). See Fig. 1. Increasing $t$ to a very large value is not informative and is not recommended in practice. On the other hand, it would be desirable to develop a data-driven method for choosing $t$, which we leave for a future work to explore.